Find the value of the variable and LN if M is between L and N.
LM = 5a, MN = 3a, LM = 40

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52879)

.

Yes, I really misread . . .

It was a regular trap . . . / as usual . . .



Answer by greenestamps(13209)



The other tutor misread the given information....

LM=5a=40, so a=40/5=8

LN=LM+MN=5a+3a=8a=8(8)=64

ANSWERS: a=8; LN=64